# Disentangling Brillouin’s negentropy law of information and Landauer’s law on data erasure

### v1

@@ Abstract
The link between information and energy introduces the observer and his knowledge into the understanding of a fundamental quantity of physics. Two approaches compete to account for this link, Brillouin’s negentropy law of information and Landauer’s law on data erasure, which are often confused. The first, based on the Clausius’ inequality and Shannon’s mathematical results is very robust, while the second, based on the simple idea that information needs a material embodiment (data-bits) is today perceived as more physical and prevails. In this paper, we show that Landauer’s idea results from a confusion between information (a global emergent concept) and data (a local material object). This confusion leads to many inconsistencies and is incompatible with thermodynamics and information theory. The reason it prevails is interpreted to be due to a frequent tendency of materialism towards reductionism, neglecting emergence and seeking to eliminate the role of the observer. A paradoxical trend given that it is often accompanied by the materialist idea that all scientific knowledge nevertheless originates from observation. Information and entropy are actually emergent quantities introduced in the theory by convention.
@@ I Introduction
For energy, there is no such definition, neither of the first type nor of the second. This is why “in physics today, we have no knowledge of what energy is” (R. Feynmann [1]). As a working definition, one can say that “in the everyday world outside of the formal language of physics, energy is the ability to do anything… Insofar as “does” means “produces a change in what was before”, energy is philosophically the determinant of all observable change”(E. Hecht) [2]). With this qualitative definition, there is no doubt that knowledge and information would be a priori forms of energy. An example is given by Maxwell [3]: an intelligent being (a demon) produces mechanical work from the information they get about a device. If energy cannot come from nowhere according to the principle of conservation, we see that, a posteriori as well, information is energy.
@@
Is that enough? No, according to Landauer. Energy is a physical quantity, information must be too [8, 9, 10]. But Landauer means by physical, something that is not abstract but tangible and material, in other word an hardware set of data-bit. Being materialized, information (understand data-bits) must be erased prior to acquisition. Hence, the Landauer’s idea: the energy cost for the acquisition of information is paid by this preliminary erasure step. This is the Landauer’s law on data-erasure.
@@
We also adopt an epistemological position throughout this article, but it will be explicit from the outset. It will be that of neo-positivists [13, 14, 15, 16] and those who influenced them [17, 18, 19, 20]. In brief: 1) only two sources of knowledge are admitted, logic and observation; 2) an explanation is a deduction of observations (past and future) from theory; 3) to fill the gap in our ignorance, the latter must include conventions chosen according to their convenience and to the economy of thought they allow. The last part of the paper shows how this position permits to analyze Landauer’s reductionist idea (beyond its inconsistencies) and how to include emergentism into materialism.
@@ II.1 Thermodynamics: principle versus law
As for the first principle, it states that the energy of the universe is conserved regardless of the change undergone by a system. In other words, the energy balance is always zero. One could imagine that the first principle also derives from precise measurements of the energy balance of various changes, which would lead to the conclusion that it is apparently always zero, thus allowing these observations to be generalized and elevated to the rank of a universal law. Of course, this assumes that we are able to recognize (identify) energy, whatever form it takes, when we see it. But that is not how it happened, because there is absolutely nothing that allows us to recognize something a priori as being a form of energy before this form has been added to the list of those known. And this is precisely this kind of measurements that allows for such an addition. The first principle does not follow from inductive reasoning. It is actually the hidden definition of what energy is, because there is no other [1].
@@ II.2 Statistical mechanics: reductionism versus emergentism
In reality, the force exerted on the piston is an emergent property, just like pressure and entropy. It cannot be derived consistently from the microscopic scale and the laws we already have at our disposal. An additional ingredient is necessary that must be postulated upstream in the theory by the conventional definition we give of macroscopic equilibrium. Just as a metal spring minimizes its potential energy at equilibrium, it would be convenient if a volume of gas would optimize something. In thermodynamics we already has the second law that says that $\mathop{}\!\mathrm{d}S\geq 0$ for an isolated system. If we postulate that entropy is maximum at equilibrium the trick is done [28] and we have the restoring force we are looking for. The equilibrium would become an attractor.
@@ III.2 Brillouin’s negentropy law of information
Let us consider a body of mass $m$, lifted by a height $\Delta h$. Classical mechanics tell us that the increase in potential energy is $mg\Delta h$. This result do not impose any particular steps for the change in height and do not specify where the mechanical work is done. However, the result concerning the net change in potential energy is general and valid in all cases, regardless of the path taken.
@@ III.3 Landauer’s law on data erasure
Landauer’s law can be examined from different angles.
The first concerns its general validity in the restricted context of data erasure. In reality, even within this context, several points of Landauer’s reasoning are questionable: the imperative to begin erasure with expansion, and the necessity of this expansion to be thermodynamically irreversible. Counterexamples have been provided [39, 40, 41] which invalidate its generality.
The second angle concerns the validity of Landauer’s law on data erasure considered as the missing link between information and energy, something that would replace Brillouin’s law of information. In the following section, we focus on the inconsistencies this idea introduces by confusing information and data.
@@ IV.1 Total versus incomplete information
In thermodynamics, setting a data bit into the S state, by removing the partition between two compartments containing a single particle (1st step of Landauer erasure), is reversible or not depending on whether the initial value of the bit is known or not [42] (see Fig. 4). This problem is the same as that of the Gibbs paradoxes and the solution also: thermodynamics considers the case where the observer has an incomplete information about the system, whereas Landauer in his reasoning implicitly assumes that they have the full information at their disposal. There is no solution to this paradox other than those that reintroduce the observer, because in reality, taking these two cases into account (as Bennett does [43], a defender of Landauer’s thesis) already amounts to taking the observer into account.
@@ IV.2 Global concept versus local object
A fundamental property of information is that it cannot be given twice. We can duplicate bits of data encoding certain information so that the storage space occupied is twice as large as before, but the corresponding amount of information remains unchanged. Conversely, if multiple copies of the same data bit are erased, the corresponding information will only be lost when the last copy is erased. To assert that information has been erased, one must take into account not only what happened locally, but also the existence or not of a copy somewhere in a more global space.
@@
The denial of the difference between information and data and the attempt to materialize information is culminating with the supposed “information-mass equivalence”[46, 12, 47, 48] that leads to paradoxes in relation with the corresponding mass deficit. For instance in the framework of this equivalence, consider a data-bit encoding a bit of information. Duplicate the bit and make the original and the copy physically independent. Erase one of the two. The quantity of information remains unchanged, so that the erased data-bit does not display any mass deficit. Erase the second bit. Information is lost that corresponds to a mass deficit. How does the second bit “knows” that the first has been erased? This contradicts the hypothesis of independence. But if we suppress this hypothesis and assert that independence is impossible, does not that mean that information is not a property of the local data bit, but a property of a global system?
Yet another variation of this paradox. Consider a data-bit, encoding for a bit of information, whose initial value is 0. Erase the data-bit (set its value to 0). There is no longer information stored by the data bit. Erase the bit again. What is the difference between the two erasures? If there is no difference, an eventual mass deficit should not be due to any loss of information. It there is a difference, how the data-bit “knows” that it has been already erased? Only a global observer could.
A data bit has a physical meaning at all scales: from that of a single particle in a two-compartment box, to the macroscopic scale. But information is meaningful only when it involves an observer, information is meaningful only at macroscale. Information, even the smallest piece, is an emergent concept.
@@ IV.3 Pair of states versus path
Consider a system changing from A to B, with $S_{\text{A}}<S_{\text{B}}$. The change is logically irreversible (the information we have decreases), thus (according to Landauer), it is also always thermodynamically irreversible. But the change in the other direction from B to A is logically reversible, thus (according to Landauer), it could be thermodynamically reversible. In summary: the change from A to B would be thermodynamically irreversible, and that from B to A would be thermodynamically reversible. What exactly would mean “thermodynamically reversible” in this context?
@@ V Conclusive epistemological point
This situation is very strange because there is absolutely no experiment, no observation, that can be explained by Landauer’s law and that would not be explained without it. On one side, the link between information and energy, as established by Maxwell or Szilard engines, is perfectly explained by Brillouin’s law. On the other side, the erasure of data bits, whatever their form, is also perfectly explained by thermodynamics. Landauer’s law adds absolutely nothing to the theory, except inconsistencies that make it incompatible with thermodynamics and information theory. So, how do we conceive its popularity? It is as if an explanation that consists solely of deducing observations (past and future) from a set of theoretical statements (principles and laws) is insufficient. Yet, that is precisely the only thing expected of theory.
The popularity of Landauer’s law can only be due to a misconception of what are concepts in physics. There is no physical theory without concepts. But concepts such as energy and information (but also that of field and many others) are not material entities, although they are fully physical, in the sense that they play a crucial role in current physical theories. They are, in fact, conventions. The meaning of any concept, what it encompasses, its definition, is a convention because we cannot see it and consequently, establish a direct link between what we see and we mean. We cannot show a concept to someone.

### v2

@@ Abstract
The link between information and energy introduces the observer and their knowledge into the understanding of a fundamental quantity of physics. Two approaches compete to account for this link, Brillouin’s negentropy law of information and Landauer’s law on data erasure, which are often confused. The first, based on the Clausius’ inequality and Shannon’s mathematical results is very robust, while the second, based on the simple idea that information needs a material embodiment (data-bits) is today perceived as more physical and prevails. In this paper, we show that Landauer’s idea results from a confusion between information (a global emergent concept) and data (a local material object). This confusion leads to many inconsistencies and is incompatible with thermodynamics and information theory. The reason it prevails is interpreted to be due to a frequent tendency of materialism towards reductionism, neglecting emergence and seeking to eliminate the role of the observer. A paradoxical trend given that it is often accompanied by the materialist idea that all scientific knowledge nevertheless originates from observation. Information and entropy are actually emergent quantities introduced in the theory by convention.
@@ I Introduction
For energy, there is no such definition, neither of the first type nor of the second. This is why “in physics today, we have no knowledge of what energy is” (R. Feynmann [1]). And this is not a joke; Feynman develops the idea over more than one page.
As a working definition, one can say that “in the everyday world outside of the formal language of physics, energy is the ability to do anything… Insofar as “does” means “produces a change in what was before”, energy is philosophically the determinant of all observable change”(E. Hecht) [2]). With this qualitative definition, there is no doubt that knowledge and information would be a priori forms of energy. An example is given by Maxwell [3]: an intelligent being (a demon) produces mechanical work from the information they get about a device. If energy cannot come from nowhere according to the principle of conservation, we see that, a posteriori as well, information is energy.
@@
Is that enough? No, according to Landauer. Energy is a physical quantity, information must be too [8, 9, 10]. But Landauer means by physical, something that is not abstract but tangible and material, in other word an hardware set of data-bits. Being materialized, information (understand data-bits) must be erased prior to acquisition. Hence, the Landauer’s idea: the energy cost for the acquisition of information is paid by this preliminary erasure step. This is the Landauer’s law on data-erasure.
@@
We also adopt an epistemological position throughout this article, but it will be explicit from the outset. It will be that of neo-positivists [13, 14, 15, 16] and those who influenced them [17, 18, 19, 20]. In brief: 1) only two sources of knowledge are admitted, logic and observation; 2) an explanation is a deduction of observations (past and future) from theory; 3) the theory includes inductive laws (generalization of observations), but it must necessarily also include conventions chosen according to their convenience and to the economy of thought they allow. The last part of the paper shows how this position permits to analyze Landauer’s reductionist idea (beyond its inconsistencies) and how to include emergentism into materialism.
@@ II.1 Thermodynamics: principle versus law
As for the first principle, it states that the energy of the universe is conserved regardless of the change undergone by a system. In other words, the energy balance is always zero. One could imagine that the first principle also derives from precise measurements of the energy balance of various changes, which would lead to the conclusion that it is apparently always zero, thus allowing these observations to be generalized and elevated to the rank of a universal law. Of course, this assumes that we are able to recognize (identify) energy, whatever form it takes, when we see it. But that is not how it happened, because there is absolutely nothing that allows us to recognize something a priori as being a form of energy before this form has been added to the list of those known. And this is precisely this kind of measurements that allows for such an addition. The first principle does not follow from inductive reasoning. It is actually the hidden, but very incomplete, definition of what energy is. Incomplete because if a quantity is conserved, meaning constant over time, any function of this quantity is also conserved and constant over time. So that, if a quantity is conserved, many others are also conserved. Among them, what should we understand by ”energy”? Despite its incomplete nature, we must be content with this definition because there is no other [1].
@@ II.2 Statistical mechanics: reductionism versus emergentism
In reality, the force exerted on the piston is an emergent property, just like pressure and entropy. It cannot be derived consistently from the microscopic scale and the laws we already have at our disposal. An additional ingredient is necessary that must be postulated upstream in the theory by the conventional definition we give of macroscopic equilibrium. Just as a metal spring minimizes its potential energy at equilibrium, it would be convenient if a volume of gas would optimize something. In thermodynamics we already have the second law that says that $\mathop{}\!\mathrm{d}S\geq 0$ for an isolated system. If we postulate that entropy is maximum at equilibrium the trick is done [28] and we have the restoring force we are looking for. The equilibrium would become an attractor.
@@ III.2 Brillouin’s negentropy law of information
Let us consider a body of mass $m$, lifted by a height $\Delta h$. Classical mechanics tell us that the increase in potential energy is $mg\Delta h$. This result does not impose any particular steps for the change in height and does not specify where the mechanical work is done. However, the result concerning the net change in potential energy is general and valid in all cases, regardless of the path taken.
@@ III.3 Landauer’s law on data erasure
Landauer’s law can be examined from two different angles:
  1. 1.
    Restricted context of data erasure: does erasing one data-bit really require at least $T\ln 2$ (Landauer’s limit) of work?
    To my knowledge, all the authors (see e.g. [39, 40, 41, 42, 43, 44]) who have addressed this question, actually used the Landauer’s procedure that consists in: a) free expansion of the phase space; b) reversible compression of the phase space. By doing this the authors simply test the 2nd law and not the new point of Landauer’s principle, so that their results are not surprising because there is absolutely no error in the calculation made by Landauer. What is new in Landauer’s assertion? It is the claim that there is no alternative erasing procedure. Thus, the only way to confirm Landauer’s principle would be to search, in vain, for an alternative. However, this is not what was done. In reality, even within the restricted context of data erasure, several points of Landauer’s reasoning are questionable: the imperative to begin erasure with expansion, and the necessity of this expansion to be thermodynamically irreversible. Counterexamples have been provided [45, 46, 47] which invalidate its generality.
  2. 2.
    Broader context of information acquisition and loss: does data erasure equate to information loss? Can Landauer’s law on data erasure be considered as the missing link between information and energy, something that would replace Brillouin’s law of information? In the following section, we focus on the inconsistencies this idea introduces by confusing information and data.
@@ IV.1 Total versus incomplete information
In thermodynamics, setting a data-bit into the S state, by removing the partition between two compartments containing a single particle (1st step of Landauer erasure), is reversible or not depending on whether the initial value of the bit is known or not [48] (see figure 4). This problem is the same as that of the Gibbs paradoxes and the solution also: thermodynamics considers the case where the observer has an incomplete information about the system, whereas Landauer in his reasoning implicitly assumes that they have the full information at their disposal. There is no solution to this paradox other than those that reintroduce the observer, because in reality, taking these two cases into account (as Bennett does [49], a defender of Landauer’s thesis) already amounts to taking the observer into account.
@@ IV.2 Global concept versus local object
A fundamental property of information is that it cannot be given twice. We can duplicate bits of data encoding certain information so that the storage space occupied is twice as large as before, but the corresponding amount of information remains unchanged. Conversely, if multiple copies of the same data-bit are erased, the corresponding information will only be lost when the last copy is erased. To assert that information has been erased, one must take into account not only what happened locally, but also the existence or not of a copy somewhere in a more global space.
@@
The denial of the difference between information and data and the attempt to materialize information is culminating with the supposed “information-mass equivalence”[50, 12, 51, 52] that leads to paradoxes in relation with the corresponding mass deficit. For instance in the framework of this equivalence, consider a data-bit encoding a bit of information. Duplicate the bit and make the original and the copy physically independent. Erase one of the two. The quantity of information remains unchanged, so that the erased data-bit does not display any mass deficit. Erase the second bit. Information is lost that corresponds to a mass deficit. How does the second bit “knows” that the first has been erased? This contradicts the hypothesis of independence. But if we suppress this hypothesis and assert that independence is impossible, does not that mean that information is not a property of the local data-bit, but a property of a global system?
Yet another variation of this paradox. Consider a data-bit, encoding for a bit of information, whose initial value is 0. Erase the data-bit (set its value to 0). There is no longer information stored by the data-bit. Erase the bit again. What is the difference between the two erasures? If there is no difference, an eventual mass deficit should not be due to any loss of information. It there is a difference, how the data-bit “knows” that it has been already erased? Only a global observer could.
A data-bit has a physical meaning at all scales: from that of a single particle in a two-compartment box, to the macroscopic scale. But information is meaningful only when it involves an observer, information is meaningful only at macroscale. Information, even the smallest piece, is an emergent concept.
@@ IV.3 Pair of states versus path
Consider a system changing from A to B, with $S_{\text{A}}<S_{\text{B}}$. The change is logically irreversible (the information we have decreases), thus (according to Landauer), it is also always thermodynamically irreversible. But the change in the other direction from B to A is logically reversible, thus (according to Landauer), it could be thermodynamically reversible. In summary: the change from A to B would be thermodynamically irreversible, and that from B to A would be thermodynamically reversible. What exactly would “thermodynamically reversible” mean in this context?
@@ V Conclusive epistemological point
This situation is very strange because there is absolutely no experiment, no observation, that can be explained by Landauer’s law and that would not be explained without it. On one side, the link between information and energy, as established by Maxwell or Szilard engines, is perfectly explained by Brillouin’s law. On the other side, the erasure of data-bits, whatever their form, is also perfectly explained by thermodynamics. Landauer’s law adds absolutely nothing to the theory, except inconsistencies that make it incompatible with thermodynamics and information theory. So, how do we conceive its popularity? It is as if an explanation that consists solely of deducing observations (past and future) from a set of theoretical statements (principles and laws) is insufficient. Yet, that is precisely the only thing expected of theory.
The popularity of Landauer’s law can only be due to a misconception of what are concepts in physics. There is no physical theory without concepts. But concepts such as energy and information (but also that of field and many others) are not material entities, although they are fully physical, in the sense that they play a crucial role in current physical theories. They are, in fact, conventions. The meaning of any concept, what it encompasses, its definition, is a convention because we cannot see it and consequently, establish a direct link between what we see and what we mean. We cannot show a concept to someone.